### v1

@@ Abstract
Distributed Hash Table Key-Value Store Surrogate Model RDMA
@@ 1 Introduction
During the last years, several research groups have shown the benefit of surrogate models for accelerating parallel simulations [18, 9, 2, 3]. While Raissi et al. [18] propose physics-informed neural networks, De Lucia et al. use decision trees to speed up geochemical simulations [3]. Reaktoro [9] and POET [2] make use of fast caches within their surrogate approach. Reaktoro uses so called On-Demand Machine Learning (ODML) to speed up chemical calculations. ODML works by extrapolating new chemical states of multiphase systems undergoing chemical reactions. Given new inputs, Reaktoro’s ODML looks up a past calculation for which the Jacobian is known and estimates the output using Taylor extrapolation [9]. POET is a coupled reactive transport simulation which uses a distributed hash table (DHT) as fast storage for simulation results. These are reused to approximate further results in subsequent time steps. An approximated lookup is used where the modeler has to specify a possible rounding for each variable. Thus, there is a tradeoff between performance and modeling accuracy. Such cache-based surrogates work fine as long as the query and retrieval times are much faster than the full physics geochemical simulations. Starting with the relevance of key-value stores in the context of big social media applications, see for example Facebook [16], researchers spent great effort to improve the performance. One of the first steps was to accelerate the communication by introducing RDMA-capable networks [7, 15, 4, 8, 1, 5, 13, 2]. All these RDMA-capable approaches have in common that they have to solve three problems:
@@
  • Presenting three distributed hash-table architectures based on the MPI one-sided communication API, including two synchronized, and one lock-free approach which uses optimistic concurrency control.
  • An evaluation of all three approaches with a synthetic benchmark testing the DHT read and write performance shows the performance advantages of the lock-free approach by a factor of up to 1,400.
  • The lock-free DHT scales very well, both for read and write requests. For 640 participating processes, 16 Million operations per second read and 15 Million operations per second write performance was observed.
@@ 3 Improving Synchronization Methods for Distributed Hash Tables with MPI
In the following, we present three different approaches to solve the data consistency problem:
  • The original MPI-DHT implementation [2] which uses MPI’s passive target synchronization, called coarse-grained locking in the following.
  • A fine-grained locking MPI-DHT which uses MPI’s atomic operations.
The decision to utilize the MPI library for the implementation of the DHT was guided by its status as the de facto standard in HPC environments, ensuring its availability to a broad user base and providing a “plug-and-play“ solution to every MPI parallelized application.
The three MPI-DHT variants are discussed in detail in the following sections.
3.1 Coarse-Grained Lock-Based MPI-DHT
@@
Figure 2: Determination of addresses of a bucket in the MPI-DHT. [2]
@@
3.2 Fine-Grained Lock-Based MPI-DHT
@@
Regarding addressing and collision handling, the coarse-grained and fine-grained locking MPI-DHTs are identical.
@@
Data Consistency: Instead of locking the entire memory window of the target rank, the fine-grained locking mechanism locks only the bucket that is currently to be accessed. One way to approach this is to define a memory window for each bucket. This naive implementation would lead to a massive overhead in terms of memory consumption, as each memory window would return a data struct with a size dependent on the MPI implementation. For Open MPI version 5.0.2, this would result in a memory overhead of 352 bytes per bucket. This is not feasible for numerous buckets. Therefore, the implementation uses self-implemented locking on 8 byte integers and relies on MPI_Compare_and_swap and MPI_Fetch_and_op operations. The implementation is adopted from the Open MPI implementation of passive-target synchronization which uses UCX atomic operations. [^1] https://github.com/open-mpi/ompi/blob/v5.0.6/ompi/mca/osc/ucx/osc_ucx_passive_target.c In order not to violate the MPI standard by performing RMA operations outside an epoch [14, p. 588], all windows are shared locked beforehand by all processes with MPI_Win_lock_all. This was inspired by the approach employed by Schuchart et al. in their work on a lock-free message queue [19].
@@ 3.3 Lock-Free MPI-DHT
4 Evaluation
@@ 4.3 Results with Synthetic Benchmarks
We varied from 1 to 5 nodes with densely mapped NUMA nodes, leading to a maximum of 640 processes. Each process provides 1 GB of memory for the DHT. The benchmark was replicated five times, and the median values of the throughput for read-only and write-only operations, respectively, along with their respective standard deviations, are displayed in Figures 3a and 3b for the uniform distribution, and in Figures 4a and 4b for the zipfian distribution.
@@ 5 Conclusion and Future Work
This paper investigates MPI-Based DHT implementations with different approaches to achieve data consistency: a coarse-grained locking DHT, a fine-grained locking DHT, and a lock-free DHT with checksum validation.
First, an evaluation of these approaches was performed using an synthetic benchmark. The lock-free DHT outperformed both synchronized DHT approaches in terms of read and write performance, with throughput improvements of up to 3 times for read operations and up to 1,400 times for write operations.
@@
In the experiments, the lock-free DHT impressed with 16 Million operations per second read and 15 Million operations per second write performance. This makes the lock-free DHT suitable as a fast data cache within HPC applications.
The benchmark code and data is published [12] and the MPI-DHT implementations are available as open-source software  [^3] A zenodo DOI will be added..

### v2

@@ Abstract
This is an extended version of the paper "A fast MPI-based Distributed Hash-Table as Surrogate Model for HPC Applications", which appeared in the Proceedings of the 25th International Conference on Computational Science (ICCS) [12]. This extended version provides more details about the approaches, implementation, and evaluation of the MPI-DHT. Additionally, it includes a comparison of the MPI-DHT with the DAOS key-value store.
@@
Distributed Hash Table Key-Value Store Surrogate Model RDMA DAOS
@@ 1 Introduction
During the last years, several research groups have shown the benefit of surrogate models for accelerating parallel simulations [20, 9, 2, 3]. While Raissi et al. [20] propose physics-informed neural networks, De Lucia et al. use decision trees to speed up geochemical simulations [3]. Reaktoro [9] and POET [2] make use of fast caches within their surrogate approach. Reaktoro uses so called On-Demand Machine Learning (ODML) to speed up chemical calculations. ODML works by extrapolating new chemical states of multiphase systems undergoing chemical reactions. Given new inputs, Reaktoro’s ODML looks up a past calculation for which the Jacobian is known and estimates the output using Taylor extrapolation [9]. POET is a coupled reactive transport simulation which uses a distributed hash table (DHT) as fast storage for simulation results. These are reused to approximate further results in subsequent time steps. An approximated lookup is used where the modeler has to specify a possible rounding for each variable. Thus, there is a tradeoff between performance and modeling accuracy. Such cache-based surrogates work fine as long as the query and retrieval times are much faster than the full physics geochemical simulations.
Starting with the relevance of key-value stores in the context of big social media applications, see for example Facebook [18], researchers spent great effort to improve the performance. One of the first steps was to accelerate the communication by introducing RDMA-capable networks [7, 17, 4, 8, 1, 5, 15, 2]. All these RDMA-capable approaches have in common that they have to solve three problems:
@@
  • A performance comparison of the MPI-based DHT (MPI-DHT) from [2] and the server-based key-value store DAOS [11]. Despite a scaling problem, the distributed architecture of the MPI-DHT performs much better than the server-based DAOS architecture.
  • Presenting two additional distributed hash-table architectures based on the MPI one-sided communication API, including a fine-grained locking DHT, and one lock-free approach which uses optimistic concurrency control.
  • An evaluation of all three DHT approaches with a synthetic benchmark testing the DHT read and write performance shows the performance advantages of the lock-free approach by a factor of up to 1,400.
  • The lock-free DHT scales very well, both for read and write requests. For 640 participating processes, 16 Million read operations per second and 15 Million write operations per second were observed.
@@ 2 Related Work
3 Comparison of MPI-DHT and the Key-Value Store DAOS
Available implementations for production use in HPC are rare. An exception are the MPI-DHT, and the DAOS implementation [^1] https://github.com/daos-stack/daos. DAOS stands for Distributed Asynchronous Object Storage and is an open-source server-based object store designed by Intel for distributed Non-Volatile Memory (NVM) technologies [11]. We conducted a comparative evaluation between the MPI-DHT [2] and the server-based approach DAOS.
Next we describe the key design features regarding addressing, data consistency and collision handling of the two systems.
3.1 Coarse-Grained Lock-Based MPI-DHT
@@
Figure 2: Determination of addresses of a bucket in the MPI-DHT. [2].
@@
3.2 DAOS
The DAOS server is responsible for determining memory addresses, ensuring the consistency of data, and handling collisions. The protocol initiates with a request message to the server. In the case of a read operation, the server is informed of the location on the client side where the data is to be written and initiates an RDMA PUT operation. In the case of a write operation, the message includes the address on the client side where the data is to be retrieved. The server then retrieves the key-value pair to be stored by a subsequent call to a RDMA GET operation. After all RMA operations have been completed, the server sends a reply message to inform the client that the requested operation is finished.
The DAOS protocol differentiates between small and large messages. For messages smaller than 18 kilobytes, the initial request message already contains the data to be written or read, thus eliminating the need for additional RMA operations on the server side.
3.3 Test bed and Benchmark Description
We conducted the evaluation on four nodes of the Turing Cluster at the University of Potsdam. Each node is equipped with two Intel Xeon E5-26540v4 processors with 12 hardware cores each, 64 GBytes of main memory, and interconnected through RoCE ConnectX-6 Dx 100 Gbit NICs.
The nodes were configured with Rocky Linux 8.8 and Mellanox OFED driver version 1.0.1. One node hosted the DAOS server, while the remaining three nodes served as clients. Since the MPI-DHT stores the values into the RAM, DAOS was also configured to use non-persistent RAM as key-value storage. For the MPI-DHT measurements, Open MPI version 4.1.5a1 was used.
We measure the operations per second (OPS) for Read and Write operations separately across varying numbers of clients. For the DAOS evaluation, we utilized the KV-API of the library to interact with the server. The key and value sizes were set to 80 bytes and 104 bytes, respectively, in accordance with the requirements of the chemical simulation described in the POET evaluation [2], which is greater than the 16 byte key and 32 byte value sizes used for the evaluation of other key-value stores [8, 4, 22]. Keys were randomly generated using a uniform distribution where every client starts with a different seed. Each client executes 100,000 write operations and then reads all 100,000 written entries. The median read and write performance of five repetitions are presented without error bars, as the maximum coefficient of variation is 3.8%.
The benchmark was scaled from 12 to 72 clients, with an increase of 12 clients per step. The processes were mapped onto a node in a dense manner, resulting in the filling of one NUMA node after another. For the distributed MPI-DHT, each client contributed 1 GByte of its memory to the DHT.
3.4 Results: Read/Write Performance
Figure 3: Throughput of DAOS and MPI-DHT for read and write operations
Obviously, the MPI-DHT outperforms DAOS in both read and write operations, as illustrated in Figure 3. The MPI-DHT reaches its peak performance with 4.12 million reads and 1.45 million writes per second when utilizing two nodes (48 processes). The throughput of DAOS remains relatively constant and reaches its maximum read rate at 60 clients with 362 thousand operations per second and its maximum write rate with 103 thousand operations per second at 72 processes.
The median latency of the MPI-DHT is between $4\,\mu s$ and $17\,\mu s$ for read access and between $13\,\mu s$ and $57\,\mu s$ for write access. The latency of DAOS is between $56\,\mu s$ and $198\,\mu s$ for read and between $157\,\mu s$ and $698\,\mu s$ for write access. The factor of improvement of the MPI-DHT compared to DAOS lies between 8.2 - 12.5 for read and 10.1 - 15.3 for write operations.
Instead of scaling with the number of clients, MPI-DHT and DAOS demonstrate both a stagnation in throughput. The poorer DAOS performance can be attributed to the central approach and indicates the server as the bottleneck of the setup. The performance of the MPI-DHT is deeper analyzed in the following.
3.5 Discussion of MPI-DHT Performance
We identified the synchronization overhead caused by the explicit locking of a target window as specified by the MPI standard as the bottleneck. To be concise, for 72 processes 80 % (DHT_write) resp. 48 % (DHT_read) of one call are spent in MPI_Win_Lock and MPI_Win_Unlock. There are two reasons for this high synchronization overhead. Firstly, the MPI-DHT treats the entire DHT memory of one process as a single window, which results in the unnecessary locking of the entire memory while only one bucket is accessed. Secondly, in the Open MPI implementation, the locking mechanism involves busy-waiting loops utilizing compare-and-swap, atomic fetch, and atomic fetch-and-add operations to facilitate exclusive and shared locking. However, this busy-waiting loop can potentially generate substantial network traffic in typical Readers & Writers scenarios where the target memory is either exclusively locked or intended to be exclusively locked, but shared locks are already acquired.
Given these limitations, we propose two new DHT synchronization approaches.
The following presents two new approaches to address the data consistency problem of the MPI-DHT:
  • A fine-grained locking MPI-DHT which uses MPI’s atomic operations.
The decision to continue using the MPI library for the implementation of the DHT was guided by its status as the de facto standard in HPC environments. This ensures the library’s availability to a broad user base and provides a “plug-and-play“ solution for every MPI-parallelized application.
The two new MPI-DHT variants are discussed in detail in the following sections.
4.1 Fine-Grained Locking MPI-DHT
@@
Regarding addressing and collision handling, the coarse-grained locking (see Section 3.1) and the fine-grained locking MPI-DHT are identical.
@@
Data Consistency: Instead of locking the entire memory window of the target rank, the fine-grained locking mechanism locks only the bucket that is currently to be accessed. One way to approach this is to define a memory window for each bucket. This naive implementation would lead to a massive overhead in terms of memory consumption, as each memory window would return a data struct with a size dependent on the MPI implementation. For Open MPI version 5.0.2, this would result in a memory overhead of 352 bytes per bucket. This is not feasible for numerous buckets. Therefore, the implementation uses self-implemented locking on 8 byte integers and relies on MPI_Compare_and_swap and MPI_Fetch_and_op operations. The implementation is adopted from the Open MPI implementation of passive-target synchronization which uses UCX atomic operations. [^2] https://github.com/open-mpi/ompi/blob/v5.0.6/ompi/mca/osc/ucx/osc_ucx_passive_target.c In order not to violate the MPI standard by performing RMA operations outside an epoch [16, p. 588], all windows are shared locked beforehand by all processes with MPI_Win_lock_all. This was inspired by the approach employed by Schuchart et al. in their work on a lock-free message queue [21].
@@ 4.2 Lock-Free MPI-DHT
5 Evaluation of the MPI-DHT with Different Synchronization Methods
@@ 5.3 Results with Synthetic Benchmarks
We varied from 1 to 5 nodes with densely mapped NUMA nodes, leading to a maximum of 640 processes. Each process provides 1 GB of memory for the DHT. The benchmark was replicated five times, and the median values of the throughput for read-only and write-only operations, respectively, along with their respective standard deviations, are displayed in Figures 4a and 4b for the uniform distribution, and in Figures 5a and 5b for the zipfian distribution.
@@ 6 Conclusion and Future Work
While server-based distributed key-value stores are more common, the setup and maintenance costs of dedicated hardware are questionable. In a preliminary study, Intel’s server-based solution DAOS was outperformed by the distributed solution MPI-DHT with coarse-grained locking. However, the MPI-DHT had scaling issues due to synchronization overhead.
Due to this fact, this paper investigates three MPI-Based DHT implementations with different approaches to achieve data consistency: a coarse-grained locking DHT, a fine-grained locking DHT, and a lock-free DHT with checksum validation.
First, an evaluation of these approaches was performed using a synthetic benchmark. The lock-free DHT outperformed both locking DHT approaches in terms of read and write performance, with throughput improvements of up to 3 times for read operations and up to 1,400 times for write operations.
@@
In the experiments, the lock-free DHT impressed with a performace of 16 Million read operations per second and 15 Million write operations per second. This makes the lock-free DHT suitable as a fast data cache within HPC applications.
The benchmark code and data are published on Zenodo [13] and the MPI-DHT implementations are available as open-source software [14].